OpenStudy (anonymous):

EASY QUESTION! DOES NOT REQIRE CALCULATION! JUST TECNICAL QUESTION! f(x) = 26(1.3)x All I need to know is what part is the slope and what part is the y-intercept.

OpenStudy (anonymous):

\[f(x)=26(1.3)^x\]Like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Y intercept is when x=0. Set x=0 and solve for y.

OpenStudy (anonymous):

(y is same as f(x) )

OpenStudy (anonymous):

\[f(x)=26(1.3)^0=?\]

OpenStudy (anonymous):

the y intercept is eather 26 or 1.3, I think its 26 but I just want to make sure...

OpenStudy (anonymous):

\[Anything^0=1\]

OpenStudy (anonymous):

So, the y-intercept is 26 .

OpenStudy (anonymous):

so then my y intercept is 26 and my slope is 1.3?

OpenStudy (vishweshshrimali5):

No, the slope of this curve is not constant.

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @student_basil \[f(x)=26(1.3)^x\]Like this ? \(\color{blue}{\text{End of Quote}}\) Just confirming, are you sure this is the equation and not this \(f(x) = 26(1.3)x\) ?

OpenStudy (vishweshshrimali5):

@potato_donuts

OpenStudy (anonymous):

im not sure now, im pretty sure its not to the power of x and just times x. because it shouldn't have a curv

OpenStudy (vishweshshrimali5):

Then, can you please check the question again from the source ?

OpenStudy (anonymous):

\[f(x)=26(1.3)x~~~~~~~~~~~~~->~~~~~~~~~~~~~~~f(x)=33.8x\]

OpenStudy (anonymous):

then y intercept is 0, and the slope is 33.8

OpenStudy (vishweshshrimali5):

I am not so sure if it is even this @student_basil . Wouldn't one directly ask for f(x) = 33.8x in that case ?

OpenStudy (vishweshshrimali5):

I think its some different f(x).

OpenStudy (anonymous):

Well, if x is not an exponent...

OpenStudy (anonymous):

im gonna close this question because im way more confused than before

OpenStudy (vishweshshrimali5):

Yeah I am also stuck there.